JET-COOLED ROTATIONAL STUDIES AT THE VALLADOLID LABORATORY

\underline{J. L. ALONSO}, J. C. LÓPEZ, A. LESARRI, S. MATA, S. BLANCO, M. E. SANZ and S.ANTOLÍNEZ, Departamento de Química Física, Facultad de Ciencias, Universidad de Valladolid, 47005 Valladolid, Spain, E-mail: jlalonso@qf.uva.es, http://www.uva.es/grupos/gem.

The low-temperature environment of a free jet expansion provides an excellent laboratory for probing the conformational behaviour of hydrogen bonding. The first observation of axial and equatorial hydrogen-bonded complexes has been reported for tetrahydropyran and pentamethylene sulphide complexes with HCl and HF.^{a} Both six-membered rings present two nonequivalent lone pairs at the O or S atoms and the formation of hydrogen bonds gives rise to two different axial and equatorial conformers. Recently, we have also reported axial and equatorial hydrogen bonds for trimethylene sulphide⋯HCl.^{b} For isolated trimethylene sulphide the ring-puckering tunnellig motion makes equivalent the nonbonding electron pairs at the S atom. This equivalence is broken by effect of complexation, giving rise to axial and equatorial conformers. In this contribution we present the results on the related pentamethylene and trimethylene sulphide⋯HF complexes. An unsubstituted alkane such as methane might act as a proton donor only in certain extraordinary circunstances. However alkanes appear capable of forming hydrogen bonds when sufficiently activated by neighboring electronegative substituents. In this context, the C-H⋯O interaction between the substituted alkane trifluoromethane and oxirane has been analyzed by MB-FTMW spectroscopy for the parent and ^{13}C isotopomers. A r_{o}-like structure has been derived and the barrier to internal rotation of the CF_{3} group has been determined from the observed A-E splittings. A MB-FTMW Spectrometer incorporating a laser ablation source has been constructed in our lab which avoids the use of high temperature techniques commonly used for vaporising low-volatility compounds. As a preliminary test, the rotational spectrum of 1,3,5-trithiane (m.p. 216 ^{o}C) has been observed. The sensitivity of the instrument proved to be high enough to observe ^{33}S isotopic species in natural abundance.

aS. Antolínez, J.C. López and J.L. Alonso, Angew. Chem. Int. Ed., 1999, 38,1772-1774. M.E. Sanz, J. C. López and J.L. Alonso, Chem. Eur. J., 1999, 5, 3293-3298. S. Antolínez, J.C. López and J.L. Alonso, ChemPhysChem, 2001, 2, 114-117.

bM.E. Sanz, A. Lesarri, J.C. López and J.L. Alonso, Angew. Chem. Int. Ed., 2001,(in press).